Single Plantwide Factory Overhead Rate Salty Sensations Snacks Company manufactures three types of snack foods: tortilla chips, potato chips, and pretzels. The company has budgeted the following costs for the upcoming period: Factory depreciation $18,717 Indirect labor 46,385 Factory electricity 5,290 Indirect materials 10,986 Selling expenses 26,041 Administrative expenses 14,648 Total costs $122,067 Factory overhead is allocated to the three products on the basis of processing hours. The products had the following production budget and processing hours per case: Budgeted Volume (Cases) Processing Hours Per Case Tortilla chips 3,900 0.12 Potato chips 2,700 0.15 Pretzels 3,600 0.10 Total 10,200 If required, round all per-case answers to the nearest cent. a. Determine the single plantwide factory overhead rate. $ per processing hour b. Use the factory overhead rate in (a) to determine the amount of total and per-case factory overhead allocated to each of the three products under generally accepted accounting principles. Total Factory Overhead Per-Case Factory Overhead Tortilla chips $ $ Potato chips Pretzels Total $ Feedback a. First calculate: Budgeted production volume for each product x Processing hours. Add together to obtain total processing hours. Recall that factory overhead does not include selling and administrative expenses. Divide the total of overhead costs by total processing hours to obtain the overhead rate. b. Calculate: Overhead Rate in Req. (a) x Processing hours for each product = Overhead for each. Add all overhead amounts to obtain the total. Divide individual overhead amounts by budgeted production volume for each product to obtain overhead per unit.
Answer #1

Estimated processing hour = (3900*0.12+2700*.15+3600*.10) = 1233 hour

a) Single plantwide factory overhead rate = (122067-14648 -26041 ) / 1233 = 66 per processing hour

b) Calcualte total amount and per case

Total Per case
Tortilla chips 3900*0.12*66 = 30888 66*.12 = 7.92
Potato chips 2700*0.15*66 = 26730 66*.15 = 9.9
Pretzels 3600*0.10*66 = 23760 66*.10 = 6.6